OpenStudy (anonymous):

Solve, using the substitution method. y = 3x + 5 4x – y = 5

OpenStudy (anonymous):

4x-3x+5=5

OpenStudy (anonymous):

X=0

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

Right? :)

hartnn (hartnn):

4x - (3x+5) = 5

hartnn (hartnn):

- (3x+5) = ... ?

OpenStudy (anonymous):

Oh... I didn't know you had to solve that too :(

OpenStudy (anonymous):

-3x-5

hartnn (hartnn):

yes, now continue

OpenStudy (anonymous):

-3x=10

hartnn (hartnn):

you mean 4x-3x = 10 ?

OpenStudy (anonymous):

Oh, yea :P X=10

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

So y=35

hartnn (hartnn):

sure it is :)

OpenStudy (anonymous):

Yesssss!! Your the best! Thanks so much

hartnn (hartnn):

you're a very good learner too :) welcome ^_^